Let 8Z be the set of all integers that are multiples of 8. Prove that 8Z has the same cardinality as 3Z, the set of all integers multiples of 3. 


1
Expert's answer
2022-01-04T15:10:23-0500

"8\mathbb{Z}=\{8x\ |\ x\in\mathbb{Z}\}" and "3\mathbb{Z}=\{3x\ |\ x\in\mathbb{Z}\}"

We need to define a bijective map from "8\mathbb{Z}" to "3\mathbb{Z}" . If there exists such map, then "8\mathbb{Z}" and "3\mathbb{Z}" have the same cardinality.

Let us define "f:\ 8\mathbb{Z}\rightarrow 3\mathbb{Z},\" "f(8x)=3x" where "x\in \mathbb{Z}."

1) "f" is one to one:

If "f(8x_1)=f(8x_2)" , then "3x_1=3x_2" . This implies that "x_1=x_2."

2) "f" is onto:

If "y\in 3\mathbb{Z}" , then "y=3z" . We need to find "x\in8\mathbb{Z}" such that "f(x)=y."

"f(x)=f(8w)=3w" and "f(x)=y=3z" .

It means that "w=z" .

So, "x\/8=y\/3" .

Therefore, for all "y\in 3\mathbb{Z}" there exists "x=8y/3\in8\mathbb{Z}" such that "f(x)=y."


Hence, "8\mathbb{Z}" has the same cardinality as "3\mathbb{Z}."